# The formation and evolution of dust in the colliding-wind binary Apep revealed by JWST

### v1

@@ 5.3 Temperature profile
The temperature profile shown in Fig. 5 closely follows a power law. We fitted a such a parametrisation to the temperature profile, defined by
@@
where $r_{0}$ was set to 10,000 au (such that $T_{0}$ is the temperature at 10,000 au). Although the temperature profiles found for the 3 features show differences just beyond uncertainties, given the potential uncertainties in photometric calibration, we do not at this stage interpret these differences to be significant. These differences are also unlikely to be due to inaccurately measured radii, as the 3D distribution of the dust is reasonably well-understood given the close fit of the geometric model (R. White et al., submitted). Assuming that the cooling law is isotropic, we fitted the power law temperature profile with an MCMC to all sample points regardless of the feature to which they correspond, finding that $\alpha=0.31\pm 0.02$ and $T_{0}=246\pm 7$ K. Note that the uncertainties on the temperature sample points, which are estimated by fitting SEDs to individual dust features, represent the most likely uncertainties only to fit these individual SEDs. These uncertainties do not take into account the fact that the temperature at similar deprojected radii should be approximately the same, despite being located at different azimuths. We therefore assumed that these uncertainties are a factor of $\zeta$ larger if a power-law temperature profile were to be enforced, and we included $\zeta$ as a free parameter when fitting the temperature profile to estimate the more likely uncertainties under this assumption. The adjusted uncertainties are shown with the thinner error bars in Fig. 5.
@@
[…]
@@
[…]
A profile shallower than the blackbody equilibrium temperature has also been assumed in prior work in the context of WR140 (P. M. Williams et al., 2009), where a power law with $\alpha=0.38$ has been assumed for amorphous carbon dust grains. While such a slope is still shallower than the value that we observationally derived, it offers a significantly improved fit than the blackbody temperature profile, suggesting that the thermal properties of dust grains in the system could be reasonably described by amorphous carbon grains within the precision achieved with these observations.
@@ 6.1 Outflow asymmetry
6.1.1 Modulation by an eccentric orbit
One possible explanation for an azimuthally asymmetric dust expansion is that the post-shock outflow velocity could vary across the binary’s orbit, which has been inferred to be eccentric (Y. Han et al., 2020; R. White et al., submitted). In WR140, extreme orbital modulation has been found to almost completely suppresses dust production near periastron (Y. Han et al., 2022). It is conceivable that even within an episode of dust formation, modulations could be manifest not only in the dust density, but also in its geometry. Under the orbital modulation scenario, the extrema in the outflow velocity likely align with periastron/apastron, the projected position angle of which is labelled in Fig. 3 based on the orbit derived in R. White et al. (submitted), which also aligns approximately with the median position angle of the dust structure as expected (P. M. Williams, 2019). This comparison shows that smaller expansion speed and shell spacing align more closely with apastron.
Variations in the wind momentum ratio could affect the opening angle of the bow shock (K. G. Gayley, 2009), however the opening angle alone cannot explain variations in dust outflow speed as it does not affect the radial displacement of the dust’s outer edge. Faster expansion of dust launched at periastron is also contrary to the expected effect if both stellar winds were subject to the same fractional suppression in outflow speed due to radiative braking, which would reduce stellar wind speeds at periastron relative to the rest of the orbit. Hydrodynamical modelling of the wind shock is required to test whether post-shock velocities could indeed be higher at smaller binary separations.
6.1.2 Contribution from orbital motion
It is possible that gravitational reflex motion near periastron could increase the outflow speed. Assuming an orbital eccentricity of 0.8 (R. White et al., submitted), an orbital period of 193 yr derived in Section 4.4, the masses of the binary components are equal and a semi-major axis of 67 au (based on the binary angular separation and position angle, Y. Han et al., 2020, and a distance of 2.4 kpc), the orbital speed at periastron in the centre-of-mass frame is approximately 40 km s-1 or 4.5 mas yr-1 at a distance of 2.4 kpc. For an unequal-mass binary, this value could increase by up to a factor of 2, which would just be comparable to the amplitude of modulation that we observe. Under this scenario, we expect the position angle of the fastest dust outflow to align with the projected velocity vector at periastron. We indeed observe that the minimum outflow speed aligns closely with the projected position angle of apastron, as shown in Fig. 3. It would therefore appear feasible for orbital motion to contribute to the asymmetric outflow observed under these estimates.
@@ 6.1.3 Interactions with the interstellar medium
Another scenario that might be proposed is asymmetries introduced by interactions with the interstellar medium. The interstellar medium would be required to be dense ($\sim 10^{4}\,\mathrm{cm}^{3}$) in such a scenario to reshape the dust motion. Under the Epstein gas drag regime, which applies to diffuse gas like in the interstellar medium and small ($\sim 0.1\,\mu$m, R. M. Lau et al., 2023) dust grains as seen in WR stars, the gas stopping time for an interstellar medium density of 100 cm-3 is 20 kyr, which is two orders of magnitude larger than the age of the dust observed, even when assuming the gas is heated to 150 K by the WR binary. In order for the gas-drag timescale to be comparable to the age of the dust, the system must be situated and moving relative to a two orders of magnitude denser region of gas than estimated above. This is disfavoured given that even if the dust grains were to survive sputtering as they move through such a dense ISM, such interactions should result in further deceleration of the dust during its expansion, which we do not observe.
@@ 6.1.4 Non-spherical wind velocity field
Although the lack of edges expanding in the sky plane in certain azimuths prevents a full characterisation in all directions, the shell spacing measurements could accommodate the possibility of two periods of oscillations per revolution, especially considering that such variations in azimuth may not be sinusoidal or symmetric. One scenario that could manifest as such a trend is if the dust expansion is spherically asymmetric, specifically with a velocity field that is prolate, which would appear circularly asymmetric when viewed at a nonzero inclination relative to a pole-on perspective. It has previously been suggested, based on measurements of the sky-plane dust expansion speed (J. R. Callingham et al., 2019; Y. Han et al., 2020) and the line-of-sight spectroscopic wind speed (J. R. Callingham et al., 2019, 2020), that the wind launched by a WR component in Apep could be anisotropic by a factor of 4 (J. R. Callingham et al., 2019; Y. Han et al., 2020). The exact azimuthal variation of the dust expansion seen in projection depends on the latitudinal distribution of the velocity field, but if the projected radial expansion varies sinusoidally, then a 24 inclination in the case of Apep (R. White et al., submitted) would result in an approximately 12% faster peak projected expansion than at minimum for an outflow that is 4 times faster in the polar direction relative to the equator, assuming that the axis of symmetry of the outflow aligns with the orbital axis. This is consistent with the level of asymmetry observed in Apep. The direction of minimum expansion speed also aligns approximately along the line of nodes (see Fig. 2), which corresponds to the two directions from the binary that would reflect a slow equatorial outflow rate under such a scenario.
@@
Although dust could form at a velocity slower than the terminal wind speed in the colliding-wind region, radiation pressure and/or secondary wind interaction downstream from the initial shock is expected to subsequently accelerate the dust (P. M. Williams et al., 2009; Y. Han et al., 2022, Monnier et al., submitted) such that the post-shock dust eventually reaches a speed comparable to the pre-shock wind speed. Indeed, the dust expansion speed observed in other colliding-wind Wolf-Rayet binaries largely agree with the wind speed (P. G. Tuthill et al., 2008; R. M. Lau et al., 2020b; P. M. Williams et al., 2009; Y. Han et al., 2022). The non-spherical outflow scenario in Apep (J. R. Callingham et al., 2019) was suggested to explain a factor of 4 discrepancy between the sky-plane expansion speed of dust (Y. Han et al., 2020) and the line-of-sight wind speed measured from spectroscopy (J. R. Callingham et al., 2020). It has been proposed that rapid rotation could support the large latitudinal wind speed gradient, implying a WR component in Apep is a Galactic long-duration gamma-ray burst progenitor. While previous estimates of the dust expansion rate have relied on sub-pixel edge detection methods to search for small shifts in dust structure and carry large uncertainties, the quadrupled time baseline enabled by the 2024 epoch over which to measure this expansion would confirm the dust expansion speed to be four times lower than the wind speed of the WN component if the distance to Apep is 2.4 kpc. Note that the distance may be further than 2.4 kpc, which would reduce this discrepancy, as we discuss in Section 6.4.
@@
For context, the half-opening angle of the shock cone is determined to be $63\pm 3^{\circ}$(R. White et al., submitted), and the wind momentum ratio $0.44\pm 0.10$ (B. Marcote et al., 2021) assuming that the shock is adiabatic. Assuming the orbit determined by R. White et al. (submitted) and distance of 2.4 kpc, the periastron separation between the binary is $12\pm 2$ au. Alternative explanations for the wind speed discrepancy have also been proposed, which include a deceleration of dust expansion following its formation in the fast wind. However, this scenario is firmly rejected by the dust structures visible in the MIRI images, which are highly regular in spacing and suggest outflow at a speed that is uniform over time. It has also been proposed that the wind speeds could be highly variable, and that the fast wind speed was measured at a time of accelerated outflow. The regular spacing of the shells and its well-preserved substructures from one episode of dust to the next also argues against wind speed variations on the timescale of hundreds of years, which would have resulted in irregularly spaced shells, with faster winds destroying the regularity of the intricate structures observed. Either a non-spherical outflow or a further distance is therefore the more plausible explanation in light of the MIRI images and the asymmetric expansion that we find.
@@ 6.2 Constraints on luminosities
The stellar radii were adjusted such that the same luminosity ratio between the three stellar components was maintained and only their total luminosity was varied. This luminosity ratio was chosen to be consistent with the near-IR interferometric constraints and the near-IR SED assembled from VLT/NACO observations in Y. Han et al. (2020), which was corrected for interstellar extinction assuming $A_{v}=11.4$ (J. R. Callingham et al., 2019). To calibrate flux density to luminosity, we used WR and O star spectra from the Potsdam WR spectral database (W. R. Hamann & G. Gräfener, 2004; A. Sander et al., 2012; R. Hainich et al., 2019), including a Galactic WC model at 63,100 K with $R_{t}=1\,\mathrm{R}_{\odot}$, a Galactic WNE model at 79,400 K with $R_{t}=1\,\mathrm{R}_{\odot}$ and a Galactic O star model at 36,000 K with log($g$)[cgs] = 3.8. We used photometry from the J, H, Ks, IB_2.24, NB_3.74 and NB_4.05 filters to constrain the flux of the northern companion, but only the J-band flux for the WR binary as its SED suggests the presence of close-in hot dust which contaminates the stellar flux at longer wavelengths (Y. Han et al., 2020). We find that a WC:WN luminosity ratio of 2.7:1 is required to reproduce the near-IR binary contrast of 5:1 (Y. Han et al., 2020), and a WR binary to O star luminosity ratio of 1.5:1 is required to reproduce their near-IR flux density ratios (Y. Han et al., 2020). Note that in the near-IR, the SED could be contaminated by free-free emission, thereby biasing the luminosity estimates to be higher than the true value. However, the effect of any free-free contamination is likely limited given the broad consistency of the near-IR SED with the XSHOOTER spectrum at visible wavelengths (J. R. Callingham et al., 2020) without incorporating a free-free emission component (Y. Han et al., 2020).
@@ 6.3 Filamentary nebula
In addition to the regular concentric dust shells, the MIRI images also suggest the presence of a diffuse filamentary nebula, which is most clearly visible in the F770W filter (see Fig. 2). The fact that the nebula becomes decreasingly visible towards longer wavelengths relative to the regularly shaped spiral dust shells suggests that the filamentary nebula exhibits either strong spectral features captured at shorter wavelengths or a significantly hotter temperature. Diffuse dust structures have been observed around other WR binaries (N. Richardson, 2025), however their spectral slopes appear to be redder that observed here, although quantitative analyses of those systems is required to confirm this comparison. We performed aperture photometry of the filamentary nebula within a region of significant emission, subtracting off background emission nearby that appear largely devoid of its presence. If the filamentary dust structure is seen in primarily thermal continuum emission, we find that reproducing its negative spectral slope between the three MIRI wavelengths would require a dust temperature of at least 350 K at a projected separation of $10^{5}$ au, which is significantly hotter than the spiral dust shell at a comparable radius even if it lies within the sky plane. The different spectral slopes between the filamentary nebula and the spiral dust shells cannot be explained by the former being significantly closer to us along the line of sight, as the interstellar dust extinction levels for Apep is only a few percent at these wavelengths.
@@
Although this probability argument alone is not sufficient to establish association, we could speculate on possible mechanisms in the scenario that the filament is being shaped or influenced by the central binary that are worthy of further investigation. Evidence of helical outflows has been observed around massive, evolved stellar environments (R. M. Lau et al., 2016), and PAH features have been suggested in the outflow of WR stars (S. V. Marchenko & A. F. J. Moffat, 2017). Furthermore, the potential rapid spin in Apep could conceivably generate magnetic fields via stellar dynamos, channelling ionised particles into polar outflows. Assuming a polar outflow four times as fast as the dust shell expansion rate and similar to the WN wind speed, and a polar axis inclination equal to that of the orbital axis at 24 (R. White et al., submitted), the dust filament in the Spitzer frame must have formed over a timescale of approximately 5,000 yr. Under such a scenario, the helical shape of the filament could reflect past orbital precessions of the binary, possibly due to a companion star detected to the north of the binary (J. R. Callingham et al., 2019) via the Kozai-Lidov mechanism (R. White et al., submitted). R. White et al. (submitted) estimated the precession timescale to be on the order $10^{5}$ yr, although significant uncertainties on the mass and orbit of the companion star could allow for faster precession timescales via the eccentric Kozai mechanism. Follow-up studies that investigate the composition and motion of the dust filament relative to Apep are required to establish the connection between the two.
Figure 6: Spitzer observations of Apep at 7.9 $\mu$m. The direction orthogonal to the line of nodes (R. White et al., submitted) is indicated with dashed lines and uncertainty range with dotted lines.

### v2

@@ 5.3 Temperature profile
The temperature profile shown in Fig. 5 closely follows a power law. We fitted such a parametrisation to the temperature profile, defined by
@@
where $r_{0}$ was set to 10,000 au (such that $T_{0}$ is the temperature at 10,000 au). Although the temperature profiles found for the 3 features show differences just beyond uncertainties, given potential biases in photometric calibration, we do not at this stage interpret these differences to be due to real physical mechanisms. Instead, they likely reflect underestimated uncertainties returned by fitting the SEDs of individual dust features. For example, E1, E2 and E3 are each affected to a different extent by outer dust structures that overlap in projection, which could lower the apparent temperature, and the azimuthally asymmetric diffraction features of the bright core, which could make the temperature appear hotter. Given the reasonably well-understood geometry of the system (White2025), uncertainties on the temperature likely dominate over uncertainties on the radius. Under the assumption that the cooling law is isotropic and that the temperature at similar deprojected radii should be approximately the same, we assumed that the temperature uncertainties fitted from individual SEDs are underestimated by a factor of $\zeta$, and we included $\zeta$ as a free parameter when fitting the temperature profile. We fitted the power law temperature profile with an MCMC to all sample points regardless of the feature to which they correspond, finding that $\alpha=0.31\pm 0.02$ and $T_{0}=246\pm 7$ K. The fitted temperature profile and the adjusted uncertainties are shown with the thinner error bars in Fig. 5. We also plotted SED models for each feature corresponding to this best-fit power-law temperature profile in Fig. 10, which show reasonable fits to the SEDs and support the assumption that the 3 features follow the same temperature profile.
@@
$\int Q_{\mathrm{abs}}(a,\lambda)B_{\nu}(\lambda,T)d\nu=\langle Q_{\mathrm{abs}}\rangle_{*}L_{*}\frac{\pi a^{2}}{16\pi r^{2}},$ (5)
@@
where $A=\int_{0}^{\infty}x^{4}/(e^{x}-1)dx\approx 24.9$. This theoretical temperature profile that scales with $r^{-0.4}$ thus deviates from the $L_{*}^{1/4}$ and $r^{-1/2}$ dependence for blackbody grains.
A profile shallower than the blackbody equilibrium temperature has also been assumed in prior work in the context of WR140 (Williams2009), where a power law with $\alpha=0.38$ has been assumed for amorphous carbon dust grains. While such a slope is still shallower than the value that we observationally derived, it offers a significantly improved fit than the blackbody temperature profile, suggesting that the thermal properties of dust grains in the system could be reasonably described by amorphous carbon grains within the precision achieved with these observations.
@@ 6.1 Outflow asymmetry
6.1.1 Wind modulation by an eccentric orbit
One possible explanation for an azimuthally asymmetric dust expansion is that the post-shock outflow velocity could vary across the binary’s orbit, which has been inferred to be eccentric (Han2020; White2025). In WR140, extreme orbital modulation has been found to almost completely suppresses dust production near periastron (Han2022). It is conceivable that even within an episode of dust formation, modulations could be manifest not only in the dust density, but also in its geometry. Under the orbital modulation scenario, the extrema in the outflow velocity likely align with periastron/apastron, the projected position angle of which is labelled in Fig. 3 based on the orbit derived in White2025, which also aligns approximately with the median position angle of the dust structure as expected (Williams2019). This comparison shows that smaller expansion speed and shell spacing align more closely with apastron.
Variations in the wind momentum ratio could affect the opening angle of the bow shock (Gayley2009), however the opening angle alone cannot explain variations in dust outflow speed as it does not affect the radial displacement of the dust’s outer edge. Faster expansion of dust launched at periastron is also contrary to the expected effect if both stellar winds were subject to the same fractional suppression in outflow speed due to radiative braking, which would reduce stellar wind speeds at periastron relative to the rest of the orbit. Hydrodynamical modelling of the wind shock is required to test whether post-shock velocities could indeed be higher at smaller binary separations.
6.1.2 Velocity contribution from orbital motion
It is possible that gravitational reflex motion near periastron could increase the outflow speed. Assuming an orbital eccentricity of 0.8 (White2025), an orbital period of 193 yr derived in Section 4.4, the masses of the binary components are equal and a semi-major axis of 67 au (based on the binary angular separation and position angle, Han2020, and a distance of 2.4 kpc), the orbital speed at periastron in the centre-of-mass frame is approximately 40 km s-1 or 4.5 mas yr-1 at a distance of 2.4 kpc. For an unequal-mass binary, this value could increase by up to a factor of 2, which would just be comparable to the amplitude of modulation that we observe. Under this scenario, we expect the position angle of the fastest dust outflow to align with the projected velocity vector at periastron. We indeed observe that the minimum outflow speed aligns closely with the projected position angle of apastron, as shown in Fig. 3. It would therefore appear feasible for orbital motion to contribute to the asymmetric outflow observed under these estimates.
@@ 6.1.3 Interactions with the interstellar medium
[…]
@@ 6.1.4 Non-spherical wind velocity field
Although the lack of edges expanding in the sky plane in certain azimuths prevents a full characterisation in all directions, the shell spacing measurements could accommodate the possibility of two periods of oscillations per revolution, especially considering that such variations in azimuth may not be sinusoidal or symmetric. One scenario that could manifest as such a trend is if the dust expansion is spherically asymmetric, specifically with a velocity field that is prolate, which would appear circularly asymmetric when viewed at a nonzero inclination relative to a pole-on perspective. It has previously been suggested, based on measurements of the sky-plane dust expansion speed (Callingham2019; Han2020) and the line-of-sight spectroscopic wind speed (Callingham2019; Callingham2020), that the wind launched by a WR component in Apep could be anisotropic by a factor of 4 (Callingham2019; Han2020). The exact azimuthal variation of the dust expansion seen in projection depends on the latitudinal distribution of the velocity field, but if the projected radial expansion varies sinusoidally, then a 24 inclination in the case of Apep (White2025) would result in an approximately 12% faster peak projected expansion than at minimum for an outflow that is 4 times faster in the polar direction relative to the equator, assuming that the axis of symmetry of the outflow aligns with the orbital axis. This is consistent with the level of asymmetry observed in Apep. The direction of minimum expansion speed also aligns approximately along the line of nodes (see Fig. 2), which corresponds to the two directions from the binary that would reflect a slow equatorial outflow rate under such a scenario.
@@
Although dust could form at a velocity slower than the terminal wind speed in the colliding-wind region, radiation pressure and/or secondary wind interaction downstream from the initial shock is expected to subsequently accelerate the dust (Williams2009; Han2022, Monnier et al., submitted) such that the post-shock dust eventually reaches a speed comparable to the pre-shock wind speed. Indeed, the dust expansion speed observed in other colliding-wind Wolf-Rayet binaries largely agree with the wind speed (Tuthill2008; Lau2020b; Williams2009; Han2022). The non-spherical outflow scenario in Apep (Callingham2019) was suggested to explain a factor of 4 discrepancy between the sky-plane expansion speed of dust (Han2020) and the line-of-sight wind speed measured from spectroscopy (Callingham2020). It has been proposed that rapid rotation could support the large latitudinal wind speed gradient, implying a WR component in Apep is a Galactic long-duration gamma-ray burst progenitor. While previous estimates of the dust expansion rate have relied on sub-pixel edge detection methods to search for small shifts in dust structure and carry large uncertainties, the quadrupled time baseline enabled by the 2024 epoch over which to measure this expansion would confirm the dust expansion speed to be four times lower than the wind speed of the WN component if the distance to Apep is 2.4 kpc. Note that the distance may be further than 2.4 kpc, which would reduce this discrepancy, as we discuss in Section 6.4.
@@
For context, the half-opening angle of the shock cone is determined to be $63\pm 3^{\circ}$(White2025), and the wind momentum ratio $0.44\pm 0.10$ (Marcote2021) assuming that the shock is adiabatic. Assuming the orbit determined by White2025 and distance of 2.4 kpc, the periastron separation between the binary is $12\pm 2$ au. Alternative explanations for the wind speed discrepancy have also been proposed, which include a deceleration of dust expansion following its formation in the fast wind. However, this scenario is firmly rejected by the dust structures visible in the MIRI images, which are highly regular in spacing and suggest outflow at a speed that is uniform over time. It has also been proposed that the wind speeds could be highly variable, and that the fast wind speed was measured at a time of accelerated outflow. The regular spacing of the shells and its well-preserved substructures from one episode of dust to the next also argues against wind speed variations on the timescale of hundreds of years, which would have resulted in irregularly spaced shells, with faster winds destroying the regularity of the intricate structures observed. Either a non-spherical outflow or a further distance is therefore the more plausible explanation in light of the MIRI images and the asymmetric expansion that we find.
@@ 6.2 Constraints on luminosities
https://gitlab.mpcdf.mpg.de/szucs/bh-mie-scat/-/tree/master
[…]
The stellar radii were adjusted such that the same luminosity ratio between the three stellar components was maintained and only their total luminosity was varied. This luminosity ratio was chosen to be consistent with the near-IR interferometric constraints and the near-IR SED assembled from VLT/NACO observations in Han2020, which was corrected for interstellar extinction assuming $A_{v}=11.4$ (Callingham2019). To calibrate flux density to luminosity, we used WR and O star spectra from the Potsdam WR spectral database (Hamann2004; Sander2012; Hainich2019), including a Galactic WC model at 63,100 K with $R_{t}=1\,\mathrm{R}_{\odot}$, a Galactic WNE model at 79,400 K with $R_{t}=1\,\mathrm{R}_{\odot}$ and a Galactic O star model at 36,000 K with log(gg)[cgs] = 3.8. We used photometry from the J, H, Ks, IB_2.24, NB_3.74 and NB_4.05 filters to constrain the flux of the northern companion, but only the J-band flux for the WR binary as its SED suggests the presence of close-in hot dust which contaminates the stellar flux at longer wavelengths (Han2020). We find that a WC:WN luminosity ratio of 2.7:1 is required to reproduce the near-IR binary contrast of 5:1 (Han2020), and a WR binary to O star luminosity ratio of 1.5:1 is required to reproduce their near-IR flux density ratios (Han2020). Note that in the near-IR, the SED could be contaminated by free-free emission, thereby biasing the luminosity estimates to be higher than the true value. However, the effect of any free-free contamination is likely limited given the broad consistency of the near-IR SED with the XSHOOTER spectrum at visible wavelengths (Callingham2020) without incorporating a free-free emission component (Han2020).
@@ 6.3 Filamentary nebula
In addition to the regular concentric dust shells, the MIRI images also suggest the presence of a diffuse filamentary nebula, which is most clearly visible in the F770W filter (see Fig. 2). The fact that the nebula becomes decreasingly visible towards longer wavelengths relative to the regularly shaped spiral dust shells suggests that the filamentary nebula exhibits either strong spectral features captured at shorter wavelengths or a significantly hotter temperature. Diffuse dust structures have been observed around other WR binaries (Richardson2025), however their spectral slopes appear to be redder that observed here, although quantitative analyses of those systems is required to confirm this comparison. We performed aperture photometry of the filamentary nebula within a region of significant emission, subtracting off background emission nearby that appear largely devoid of its presence. If the filamentary dust structure is seen in primarily thermal continuum emission, we find that reproducing its negative spectral slope between the three MIRI wavelengths would require a dust temperature of at least 350 K at a projected separation of $10^{5}$ au, which is significantly hotter than the spiral dust shell at a comparable radius even if it lies within the sky plane. The different spectral slopes between the filamentary nebula and the spiral dust shells cannot be explained by the former being significantly closer to us along the line of sight, as the interstellar dust extinction levels for Apep is only a few percent at these wavelengths.
@@
Although this probability argument alone is not sufficient to establish association, we could speculate on possible mechanisms in the scenario that the filament is being shaped or influenced by the central binary that are worthy of further investigation. Evidence of helical outflows has been observed around massive, evolved stellar environments (Lau2016), and PAH features have been suggested in the outflow of WR stars (Marchenko2017). Furthermore, the potential rapid spin in Apep could conceivably generate magnetic fields via stellar dynamos, channelling ionised particles into polar outflows. Assuming a polar outflow four times as fast as the dust shell expansion rate and similar to the WN wind speed, and a polar axis inclination equal to that of the orbital axis at 24 (White2025), the dust filament in the Spitzer frame must have formed over a timescale of approximately 5,000 yr. Under such a scenario, the helical shape of the filament could reflect past orbital precessions of the binary, possibly due to a companion star detected to the north of the binary (Callingham2019) via the Kozai-Lidov mechanism (White2025). White2025 estimated the precession timescale to be on the order $10^{5}$ yr, although significant uncertainties on the mass and orbit of the companion star could allow for faster precession timescales via the eccentric Kozai mechanism. Follow-up studies that investigate the composition and motion of the dust filament relative to Apep are required to establish the connection between the two.
[…]